OpenStudy (anonymous):

Can someone explain the whole concept of solving systems by graphing? I GIVE MEDALS!!!!!!!

OpenStudy (anonymous):

Some problems can be represented by lines. The problems that are solvable by systems, when graphed, the 2 lines will intersect at one point, that one point is the answer to the problem.

OpenStudy (anonymous):

ok thank you but i meant the formulas for how to turn standard form into slope intercept form

OpenStudy (marigirl):

lets take 4x-y=5 as an example. it is in standard form at the moment when we change it to y=mx+c form ... y=4x-5 it is now in slope and intercept form. the slope is 4 (gradient) and y intercept is at -5

OpenStudy (marigirl):

changing equations in standard form to y=mx+c form (gradient and y intercept form) makes it quicker to graph.

OpenStudy (marigirl):

thanx for the medal :D

OpenStudy (anonymous):

@marigirl Good job!